OpenStudy (anonymous):

Eliminate the parameter, t, and write a rectangular equation for x =t^2 +3 and y = 2t.

OpenStudy (anonymous):

\[x = \frac{y^2}{4} + 3 \]

OpenStudy (anonymous):

I eliminated t

OpenStudy (anonymous):

I killed t ... I am math serial killer

OpenStudy (anonymous):

i got t = \[\sqrt{3-x}\]

OpenStudy (anonymous):

you said eliminate t not y

OpenStudy (anonymous):

yeah to x = \[t ^{2} + 3\] then t\[t ^{2} = 3-x\] tehn t = \[\sqrt{3-x}\]

OpenStudy (anonymous):

How do you write a rectangualr equation??

OpenStudy (anonymous):

ah I really don't know what is rectangular equation ...

OpenStudy (anonymous):

oh ok...but then to find y since y=2t what is \[2 (\sqrt{3-x}\]

OpenStudy (anonymous):

maybe it means parabola's equation

OpenStudy (anonymous):

how do i distribute the 2 to the radical?

OpenStudy (anonymous):

use my equations to find y x = y^2/4 + 3 4x = y^2 + 12 \[\sqrt{4x - 12 } =y \]